OpenStudy (gabylovesyou):

Identify the correct x-intercept and y-intercept of the graph of the equation 3x + 9y = 18. x-intercept = (6, 0); y-intercept = (0, 2) intercept = (-6, 0); y-intercept = (0, -2) intercept = (2, 0); y-intercept = (0,1/3 ) intercept = (1/3, 0); y-intercept = (0, 6)

OpenStudy (anonymous):

so you're line is 3x + 9y = 18, Set x=0 to find the y intercept, set y=0 to find the x intercept when x=0....3(0) + 9y = 18, y = 2 when y=0....3x + 9(0) = 18, x = 6 y intercept is (0,2), x intercept is (6,). So that is the first one.